I'm not sure how to simplify this problem:
x(squared) + 3
over
x(squared) + 4
is equal to 2 (radical)2. I tried multiplying the faction by its reciprocal, but I can't figure out what to do with the radical on the right side.
Thank you!

Answer by lwsshak3(11628)

I'm not sure how to simplify this problem:
x(squared) + 3
over
x(squared) + 4
is equal to 2 (radical)2. I tried multiplying the faction by its reciprocal, but I can't figure out what to do with the radical on the right side.
**
(x^2+3)/(x^2+4)=2√2
square both sides
(x^4+6x^2+9)/(x^4+8x^2+16)=4*2=8
(x^4+6x^2+9)=(8x^4+64x^2+128)
combine like terms, set equation to 0
7x^4+58x^2+119=0
let u=x^2
7u^2+58u+119
solve for u by quadratic equation
then solve for x (4 roots)